OpenStudy (anonymous):

d/dx [3e^x + 5ln x] = what are the first steps to solve?

OpenStudy (anonymous):

do you know what the derivative of \(e^x\) is?

OpenStudy (anonymous):

e^x...right?

OpenStudy (anonymous):

yes, so the derivative of \(3e^x\) is also \(3e^x\)

OpenStudy (anonymous):

do you know what the derivative of \(\ln(x)\) is ?

OpenStudy (anonymous):

Ohh ok. Thank you. What next?

OpenStudy (anonymous):

just x...right?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

\[\frac{d}{dx}[\ln(x)]=\frac{1}{x}\] right

OpenStudy (anonymous):

and so the derivative of \(5\ln(x)\) is \(\frac{5}{x}\)

OpenStudy (anonymous):

stick a plus sign between them and you are done

OpenStudy (anonymous):

ohh ok. so it is 3e^x + (5/x)

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

yes, that is it, not too hard right?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

no. thank you for your help